OpenStudy (anonymous):

3^5x+7 = 9 solve for x

OpenStudy (gorv):

subtract 7 from both side

OpenStudy (gorv):

3^5x=9-7=2 3^5x=2

OpenStudy (gorv):

now take log both side

OpenStudy (gorv):

log(3^5x)=log2 5x*log3=log2

OpenStudy (anonymous):

is it the same using natural log?

OpenStudy (anonymous):

sorry forgot to include that in problem

OpenStudy (gorv):

yeah naturallog it is

OpenStudy (anonymous):

thanks. how does the answer get to -1? I have -1 as the answer on my worksheet